On January 1, Vermont Corporation had 35,300 shares of $12 par value common stock issued and outstanding. All 35,300 shares had been issued in a prior period at $22 per share. On February 1, Vermont purchased 1,140 shares of treasury stock for $25 per share and later sold the treasury shares for $20 per share on March 1.

The journal entry to record the purchase of the treasury shares on February 1 would include a

a.credit to Treasury Stock for $28,500

b.debit to Treasury Stock for $28,500

c.credit to a gain account for $3,420

d.debit to a loss account for $3,420

A corporation has 49,347 shares of $38 par stock outstanding that has a current market value of $350 per share. If the corporation issues a 4-for-1 stock split, the market value of the stock will fall to approximately

a.$312.00

b.$87.50

c.$9.50

d.$1,400.00
Answer #1

1. The entry would be

Cash (1140*20) 22,800
Paid in capital in excess of par 5,700
Treasury stock (1140*25) 28,500

Option A

2.

Outstanding shares after split = 49,347*4/1 = 197,388

Market value after split = (49,347*38)/197,388 = 9.50

Option C